# CEval: A Benchmark for Evaluating Counterfactual Text Generation

### v1

@@ 6 Conclusion
We propose CEval to raise awareness and suggest a standard way to evaluate counterfactual generation in text data, an area gaining traction in the NLP community. Given the nature of textual data, we emphasize the importance of both, counterfactual metrics and text quality. Our benchmark not only fosters standardized comparisons, but also analyzing strengths and weaknesses of individual methods. Our instantiation of the benchmark with representative methods revealed that counterfactual methods excel in counterfactual metrics, but suffer degraded text quality. In turn, LLMs produce high quality text, but struggle to reliably flip the label. Combining the strengths of both approaches opens avenues for future research, e.g., by guiding an LLM’s output with supervision from the target classifier. Low diversity within methods and high diversity across methods highlights the need to conduct robustness analyses of the target classifier with multiple methods. An imperfect flip rate of human generated text (and partly LLMs) can reveal predictive performance issues of the target classifier. The target classifier in our experiments seems to be poorly calibrated and we plan to investigate the impact of well calibrated classifiers in future work. By providing CEval as an open-source Python library, we encourage the community to contribute additional methods and ensuring future work follows the same standards.
@@
We employ default hyperparameters for each method and straightforward prompts with LLMs, which may not be optimal for the task at hand and could be furtrher improved by hyperparameter optimization and prompt engineering.
This benchmark solely evaluates the quality of counterfactual text. Further research is required to evaluate the performance of this text in downstream tasks such as classification, data augmentation, or robustness improvement. Additionally, we evaluate the metrics with a single classifier. While this classifier achieves state-of-the-art classification accuracy, our results indicate that it might not be well calibrated. Estimating to which extent our findings can be generalized requires a combination of multiple diverse classifiers in the benchmark and the application in downstream tasks.
Appendix A Method Selection Criteria
Method Type Classifier Access Reproducible code Problem Agnostic
@@ Appendix A Method Selection Criteria
Table 5: Comparison of Methods. Those methods with different types meeting all criteria (highlighted in bold) have been selected for inclusion in the benchmark.
Appendix B Generated Text Comparison Example
Method Text Predicted label
@@ Appendix B Generated Text Comparison Example
Table 6: Example that most methods failed to flip the label

### v2

@@ 7 Conclusion
We propose CEval to standardize the evaluation of counterfactual text generation, emphasizing the importance of both counterfactual metrics and text quality. Our benchmark facilitates standardized comparisons and analyzes the strengths and weaknesses of individual methods. Initial results show that counterfactual methods excel in counterfactual metrics but produce lower-quality text, while LLMs generate high-quality text but struggle to reliably flip labels. Combining these approaches could guide future research, such as using target classifier supervision to enhance LLM outputs. The diversity in method performance highlights the need for robustness analyses of target classifiers with multiple methods. Our findings also suggest that the target classifier may be poorly calibrated, warranting further investigation. Finally, we demonstrate that text quality evaluation using LLMs is robust to temperature changes. Additionally, we show that open-source LLMs, like Mistral, can serve as alternatives to closed-source models, such as ChatGPT, for evaluating text quality, thereby overcoming weaknesses of closed-source models, such as API deprecation or high costs. This leads to CEval being a fully open-source Python library, encouraging the community to contribute additional methods and to ensure that future work follows the same standards. For future work, we plan to integrate LLMs specifically designed for evaluation, such as Prometheus Kim et al. (2023), as an option for assessing text quality. Furthermore, instead of only considering the difference between instances to measure diversity, the diversity metric can be expanded to incorporate the particular types of changes, such as negation and word replacements.
@@
We employ default hyperparameters for each method and straightforward prompts with LLMs, which may not be optimal for the task at hand and could be further improved by hyperparameter optimization and prompt engineering.
This benchmark solely evaluates the quality of counterfactual text for explanation tasks. Further research is required to evaluate the performance of this text in other downstream tasks such as data augmentation with counterfactual examples or improving the robustness of the model using counterfactual examples. Additionally, we evaluate the metrics with a single BERT-based classifier. While this classifier achieves state-of-the-art classification accuracy, our results indicate that it might not be well calibrated. Estimating to which extent our findings can be generalized requires a combination of multiple diverse classifiers in the benchmark and the application in downstream tasks.
A potential exposure of ChatGPT or Mistral to the human counterfactual dataset is unlikely to impact our results, as we used these models only for evaluating text quality rather than counterfactual generation. The exposure of LLAMA-2 to human counterfactuals remains uncertain. If such exposure occurred, it could potentially influence our results for LLAMA-2, as it would help to generate better (human-like) counterfactuals. However, Fig. 3 shows a considerable distance between human-generated and LLAMA-generated counterfactuals, suggesting a low likelihood of such influence.
Ethics Statement
We use the publicly available datasets IMDB and SNLI, and employ the benchmark to evaluate existing counterfactual generation methods. None of these methods declared any ethical concerns. While the benchmark is designed to evaluate counterfactual generation methods to advance research in explainable AI, it could be misused to select the best counterfactual methods for generating potentially harmful content. One such harmful application scenario could be the generation of counterfactuals to evade a fake news detector. However, if such evasion would actually be possible without a drastic change of the semantics, the major risk stems from the counterfactual generation methods rather than from their benchmark comparison.
We strongly believe that a benchmark evaluation should be as open, fair, transparent and reproducible as possible. Therefore, we make all our source code (including benchmark evaluation and method implementation) publicly available1 and include the option to evaluate text quality metrics with the open-source LLM Mistral-7B (cf. Section 6).
Appendix A Generated Text Comparison Example
Appendix B Method Selection Criteria
Method Type Classifier Access Reproducible code Problem Agnosticity
@@ Appendix B Method Selection Criteria
Table 4: Comparison of Methods. Methods of different types that meet all inclusion criteria are highlighted in bold and are included in the benchmark.
Appendix C Correlation of Mistral and ChatGPT
Temperature 0.2 1.0
Grammar 1.0 0.89
Cohesiveness 0.94 0.89
Fluency 1.0 0.94
Table 5: Spearman correlation of method rankings assigned by the LLM models Mistral and ChatGPT across different temperature settings, demonstrating very strong correlation.
Appendix D Effect of Temperature
We evaluate the effect of temperature on the counterfactual generation process and text quality. Table 6 shows the results of LLAMA-2 with three different temperatures: 0.2, 0.6, and 1.0. Lower temperatures imply a higher likelihood of selecting the most frequent tokens and a lower likelihood of selecting less frequent tokens. Consequently, diversity is low at lower temperatures and high at higher temperatures. Perplexity is also correlated with temperature, while other metrics do not show a clear correlation. On the other hand, Figures 4 and 5 show the correlations between the same model at different temperatures, as well as the correlations between different models across various metrics. We observe a very strong correlation within the same model and a moderate correlation when using different models, suggesting that the evaluation is robust with respect to temperature.
IMDB SNLI
0.2 0.6 1.0 0.2 0.6 1.0
CF Metrics Flip Rate $\uparrow$ 0.68 0.65 0.70 0.38 0.40 0.39
$\Delta$Probability $\uparrow$ 0.67 0.66 0.69 0.32 0.33 0.33
Perplexity $\downarrow$ 40.6 39.1 41.3 54.9 55.2 57.0
Distance $\downarrow$ 50.7 48.9 58.0 4.36 4.48 4.78
Diversity $\uparrow$ 28.3 44.4 61.6 - - -
Text Quality Grammar $\uparrow$ 3.20 3.18 3.18 3.76 3.77 3.68
Cohesiveness $\uparrow$ 3.14 3.15 3.12 3.71 3.69 3.61
Fluency $\uparrow$ 3.12 3.11 3.13 3.66 3.71 3.59
Average $\uparrow$ 3.15 3.15 3.14 3.71 3.72 3.63
Table 6: Comparison of LLAMA-2 counterfactual generation with different temperatures (0.2, 0.6, and 1.0). Temperature primarily affects diversity, with minimal impact on other metrics.
LLAMA-2 MICE GBDA CREST Crowd
E N C E N C E N C E N C E N C
Grammar Cohesiveness Fluency
GPT Mistral GPT Mistral GPT Mistral
0.2 1.0 0.2 1.0 0.2 1.0 0.2 1.0 0.2 1.0 0.2 1.0
Crowd 3.58 3.56 4.62 4.61 3.60 3.53 3.77 3.73 3.56 3.51 4.48 4.43
Crest 2.71 2.66 3.71 3.73 2.74 2.72 3.03 3.00 2.70 2.66 3.88 3.82
GBDA 2.29 2.31 3.27 3.22 2.03 2.08 2.10 2.20 2.17 2.16 3.37 3.31
Mice 3.33 3.32 4.44 4.39 3.31 3.31 3.50 3.46 3.33 3.34 4.38 4.29
LLAMA-2 3.68 3.66 4.63 4.60 3.61 3.55 3.64 3.63 3.59 3.58 4.44 4.36
Table 8: Comparison of text quality evaluation using Mistral and ChatGPT (GPT-3.5 Turbo) with different temperatures (0.2 and 1.0) on SNLI dataset.
Figure 4: Pearson correlation between Mistral and ChatGPT in text quality evaluation with different temperatures (0.2 and 1.0) on the IMDB dataset. The same model with the different temperatures exhibits a strong correlation, meanwhile different models show a moderate correlation in evaluating text quality for counterfactual generation.
Figure 5: Pearson correlation between Mistral and ChatGPT in text quality evaluation with different temperatures (0.2 and 1.0) on the SNLI dataset. Text quality evaluation results of the same model with the different temperatures are strongly correlated; results from different models are moderately correlated.
Method Text Predicted
Label
@@ E N C E N C E N C E N C E N C
Table 9: Example for which most methods failed to flip the label